### v2

# BadSFL: Backdoor Attack against Scaffold Federated Learning

Abstract
Federated learning (FL) enables the training of deep learning models on distributed clients to preserve data privacy. However, this learning paradigm is vulnerable to backdoor attacks, where malicious clients can upload poisoned local models to embed backdoors into the global model, leading to attacker-desired predictions. Existing backdoor attacks mainly focus on FL with independently and identically distributed (IID) scenarios, while real-world FL training data are typically non-IID. Current strategies for non-IID backdoor attacks suffer from limitations in maintaining effectiveness and durability. To address these challenges, we propose a novel backdoor attack method, BadSFL, specifically designed for the FL framework using the scaffold aggregation algorithm in non-IID settings. BadSFL leverages a Generative Adversarial Network (GAN) based on the global model to complement the training set, achieving high accuracy on both backdoor and benign samples. It utilizes a specific feature as the backdoor trigger to ensure stealthiness, and exploits the Scaffold’s control variate to predict the global model’s convergence direction, ensuring the backdoor’s persistence. Extensive experiments on three benchmark datasets demonstrate the high effectiveness, stealthiness, and durability of BadSFL. Notably, our attack remains effective over 60 rounds in the global model and up to 3 times longer than existing baseline attacks after stopping the injection of malicious updates.
Introduction
The distribution of training data across clients plays a crucial role in the performance of Federated Learning (FL) systems. Two scenarios typically arise: 1) the training data is IID among clients; 2) A more realistic scenario where the data is non-IID, meaning the data characteristics vary significantly across clients. For IID scenarios, FedAvg (McMahan et al. 2017) stands out as the leading FL method, setting the standard for server-side model updates by aggregating model parameters from clients effectively. However, its performance deteriorates in non-IID scenarios, where data heterogeneity causes update drifts from individual clients, ultimately degrading convergence. (Li et al. 2019). To address this challenge, Scaffold (Karimireddy et al. 2020) has been proposed as a promising solution, using a control variate to correct the drift induced by the data heterogeneity.
However, while FL’s distributed training paradigm allows clients to have full control over their data for privacy protection, it also introduces a non-negligible security threat: backdoor attacks, where a malicious client can manipulate the local training process to embed targeted misbehaviors into the global model through poisoned updates (Chen et al. 2017).
Existing federated learning backdoor attacks (Bagdasaryan et al. 2020) (Xie et al. 2019) (Sun et al. 2019) (Wang et al. 2020a) (Dai and Li 2023) mainly focus on IID scenarios, where attackers compromise clients with full knowledge of the dataset distribution to craft highly effective attacks. Unfortunately, these methods have not demonstrated their feasibility and effectiveness in non-IID scenarios. Specifically, attackers lacking full knowledge of the dataset distribution often result in training backdoor models diverging significantly from the global model, severely compromising overall model performance. Consequently, directly applying attack methods from IID scenarios to non-IID scenarios is ineffective. Some efforts have performed backdoor attacks on non-IID scenarios (Alam, Sarkar, and Maniatakos 2023) (Jeong, Hwang, and Chung 2021) (Naseri, Han, and De Cristofaro 2023) (Ye et al. 2024). However, their effectiveness remains limited against advanced defenses like Scaffold, since they do not take the control variate into consideration.
In this paper, we propose BadSFL, a novel backdoor attack specifically targeting Scaffold Federated Learning (SFL), to successfully plant a backdoor function into a global model without catastrophically corrupt model performance on benign sample inference. Firstly, as the attacker only has partial knowledge of the dataset distribution in the FL system, we leverage a GAN to generate fake samples that belong to other clients to supplement the dataset, simulating a full knowledge of the dataset distribution. With a supplemented dataset for backdoor training, we get a backdoor model achieving high accuracy in both backdoor tasks and benign tasks. Secondly, we use a distinctive feature of a category as the backdoor trigger to maintain the stealthiness of BadSFL. Thirdly, we exploit the global control variate, a specific feature of SFL, as a reference to predict the global model’s convergence direction. This optimization enhances the durability of the embedded backdoor function within the global model. We evaluate the effectiveness of BadSFL on the MNIST, CIFAR-10, and CIFAR-100, demonstrating high accuracy on both backdoor and primary tasks. Moreover, the embedded backdoor function persists in the global model for over 60 rounds and lasts for 3 times longer than baseline attacks after the attacker stops injecting malicious updates. Finally, we show that BadSFL remains highly effective when using three defense methods simultaneously.
The main contributions are as follows:
  • We propose BadSFL, the first backdoor attack against scaffold federated learning on non-IID scenarios.
  • We enhance the durability of the backdoor, ensuring it persists for over 60 rounds and lasts 3 times longer than baseline attacks.
  • We conduct extensive experiments on three benchmark datasets. The results demonstrate that our method achieves high effectiveness.
Background and Related Work
Non-IID Scenarios in FL
In FL, non-IID refers to significant differences in data distributions among clients (Kairouz et al. 2021) (Li et al. 2021b) (Mendieta et al. 2022). This discrepancy between local data distributions in NON-IID scenarios can lead to inconsistencies between the local optima and the global optima. This inconsistency results in a drift in local model updates, where local models move towards their own local optima which can be far from the global optima (Karimireddy et al. 2020). Consequently, even when averaging these local models may yield a global model far from the true global optima (Li et al. 2020) (Wang et al. 2020b) (Wang et al. 2020c) (Karimireddy et al. 2020) (Lee et al. 2022), especially with numerous local epochs. Figure 1 illustrates this issue: while global optima align with local optima in IID scenarios. non-IID can cause the global optima to be distant from individual local optima, which is known as the client-drift phenomenon, leading to slow and unstable convergence in the FL training process.
Figure 1: Model Avg under IID and non-IID scenarios.
SCAFFOLD. Several FL algorithms have been proposed to address the above challenges, with SCAFFOLD (Karimireddy et al. 2020) being the most practical solution. It tackles the client-drift problem through the control variates (variance reduction techniques) for both the server and the clients. These control variates estimate the update direction of the global model and local client models and serve to correct local updates based on the drift, thereby mitigating the divergence between local and global optima. Therefore, we mainly focus on designing backdoor attacks targeting scaffold federated learning (SFL) in this paper.
Backdoor Attacks in FL
Backdoor attacks pose a significant threat to the security of deep learning models, where malicious clients embed hidden triggers within models that cause misclassification during inference while maintaining normal performance on clean data (Li et al. 2022) (Gu et al. 2019). In FL, adversaries can deploy backdoor attacks by exploiting compromised clients to manipulate local updates, thereby generating poisoned models that corrupt the global model upon aggregation (Bagdasaryan et al. 2020)(Sun et al. 2019)(Wang et al. 2020a)(Xie et al. 2019)(Zhang et al. 2022) (Li, He, and Song 2021). Backdoor attacks typically involve a combination of model replacement and data poisoning. In model replacement, the adversary substitutes legitimate models with manipulated ones (Bagdasaryan et al. 2020), while data poisoning involves injecting poisoned data, containing the backdoor trigger, into the training datasets of compromised clients. To inject these triggers, the adversary can manipulate the dataset set by flipping data labels or adding a unique pixel pattern to images (Bagdasaryan et al. 2020)(Wang et al. 2020a)(Alam, Lamri, and Maniatakos 2023). Afterward, they strategically adjust training parameters and scale updates to optimize the impact of the attack while evading detection by the anomaly detector deployed at the central aggregation server (Blanchard et al. 2017)(Nguyen et al. 2022). In this paper, we mainly focus on data poisoning attacks.
Threat Model
Attack scenarios. We consider an attacker who aims to inject backdoors into SFL to make the victim clients predict the desired wrong output over a triggered input. The attacker has partial knowledge of the full dataset in the training stage. Specifically, the attacker trains a local model with a backdoor trigger function and submits poisoned local updates along with the control variate to the server for Scaffold aggregation. During inference, the attack manipulates predictions to produce the attacker’s desired outputs when inputs from benign clients meet the trigger conditions.
Attack goal. The attack goal can be summarized as follows:
  • $\bullet$
    Effectiveness: The attacker must ensure that the backdoor function does not compromise the global model’s performance on primary tasks, maintaining high accuracy for both backdoor and benign predictions.
  • $\bullet$
    Stealthiness: the backdoor should be robust against potential defenses.
  • $\bullet$
    Durability: The backdoor should remain effective in the global model for as long as possible, even after the attacker stops participation in the training process, thereby maximizing the longevity of the attack.
Attack capability. We assume the attacker can compromise at least one client, acting as a benign client during the training process, thereby allowing him to operate covertly within the system. Additionally, participation in FL provides the attacker with full knowledge of the model structure, facilitated by the consensus among all clients on a common learning objective. This enables the insertion of backdoor triggers, modification of sample labels, and manipulation of local training updates. It is essential to note that the attacker cannot control the server or directly manipulate the aggregation procedure or the global model. They also lack access to data and models from non-compromised clients.
Challenges with Backdoor Attacks in SFL
Performing a backdoor attack in SFL presents the following challenges.
Limited knowledge: In non-IID scenarios, a primary challenge arises from the attacker’s lack of knowledge of the dataset distribution across clients. Unlike IID scenarios where a centralized understanding of the dataset facilitated manipulation, non-IID scenarios involve decentralized and diverse data distributions. This results in three issues: (1) Direct backdoor strategies can cause significant performance degradation on benign samples, leading to the rejection of the global model. (2) The variability in data distributions increases the difference between local and global models, making malicious models more detectable. (3) Averaging poisoned models with the global model degrades its performance on the primary task, as shown in Figure 2, where accuracy drops significantly when a poisoned model is aggregated.
Figure 2: Left: Primary task accuracy crushed by simple attack; Right: Backdoor accuracy dropped after stop attacking.
Control variate: In SFL, control variate (denoted as $c_{i}$) is used to correct the client drift and align local models with the global model. If attackers strictly follow protocols and use the $c_{i}$) to correct their malicious models during the triggering planting process, the effectiveness of the attack can be reduced. Conversely, if an attacker chooses to manipulate the $c_{i}$ inappropriately, introducing a malicious $c$ to the server, it could lead to a potential corruption of the global model.
Backdoor catastrophic forgetting: Catastrophic forgetting (Kirkpatrick et al. 2017) occurs when neural networks forget previously learned tasks upon learning new ones. This can cause backdoors to lose effectiveness over time, as demonstrated by Wang et al. (Wang et al. 2020a). If attackers stop uploading malicious updates, the backdoor function may eventually be erased by benign updates. As shown in Figure 2, the accuracy of backdoor tasks declines sharply over time, with the backdoor function vanishing around round 65. Although various methods (Zhang et al. 2022)(Alam, Sarkar, and Maniatakos 2023)(Dai and Li 2023)(Wen et al. 2022) have been proposed to address this issue, none have proven effective in SFL.
Algorithm 1 BadSFL attack Algorithm
Required: local datasets $D^{i}$, global model $w_{g}$, global control variate $c$, number of local epochs $E$, local learning rate $\eta_{l}$, Generator $G$, Discriminator $D$
   Update local model with $w_{p}\leftarrow w_{g}$
   Initialize Discriminator $D\leftarrow w_{g}$
   do:
      Run G for generating fake samples
      Evaluate fake sample on D
      Update G using D
   until G converges to generate target samples
   G generates samples into dataset $D_{f}$
   $D_{c}\leftarrow D_{f}+D^{i}$
   Select backdoor samples from $D_{c}$ and assign them wrong label as $D_{b}$
   $D_{p}\leftarrow D_{c}+D_{b}$
   for each epoch e = 1, …, $E$ do:
       $w_{p}=\mathop{argmin}_{w_{p}}[L(D_{p},w_{p})+L(D_{p},P_{j}(w_{p},c))].$
   end for
   $\Delta w_{p}=w_{p}-w_{g}$
   $\Delta c_{p}=\frac{1}{K*\eta_{l}}*(w_{g}-w_{p})-c$
return ($\Delta w_{p},\Delta c_{p}$)
BadSFL
Overview. To overcome the challenges, we propose BadSFL, which is detailed in Algorithm 1. BadSFL mainly consists of 4 steps: Step 1: Initialization. The attacker initiates the attack by downloading the global model $w_{g}$ and controlling the variate $c$ from the server. Subsequently, the attacker updates the local model $w_{p}$ and discriminator $D$ using the downloaded global model $w_{g}$. Step 2: GAN-based Training for Data Supplementation. The attacker performs GAN training on generator $G$ and discriminator $D$. The training terminates upon the convergence of the generator, signifying its ability to generate realistic fake samples in class $C$ that do not belong to $D^{i}$ but rather originate from other clients’ datasets. Then the generator $G$ is utilized to generate a number of samples in class $c$ forming in dataset $D_{f}$. This dataset $D_{f}$ is then merged with the attacker’s original dataset $D^{i}$ to create a new dataset $D_{c}$. Step 3: Backdoor Sample Selection and Trigger Injection. With dataset $D_{c}$, the attacker selects specific samples with a characteristic feature to serve as backdoor samples. These samples are then relabeled to a target class $x$ as the backdoor target class that is different from their original labels. The attacker organizes these manipulated samples into a separate dataset $D_{b}$ and merges it with $D_{c}$ to finalize the dataset $D_{p}$ for backdoor training. Step 4: Backdoor Model Training and Optimization. The attacker proceed to trains the local model $w_{p}$ based on the dataset $D_{p}$. During the training process, the attacker follows the equation 3 to optimize the backdoor objective. Upon convergence, the backdoor model update $\Delta w_{p}$ and the corresponding control variate $\Delta c_{p}$ are obtained and can be uploaded to the server.
GAN-based dataset supplementation. In non-IID data scenarios, directly injecting backdoor samples into dataset $D^{i}$ for training often leads to a more biased model, deviating significantly from global optima (Psychogyios et al. 2023) (Lai et al. 2022). To mitigate this, inspired by Zhang et al. (Zhang et al. 2019), the attacker can employ GAN to generate synthetic samples that resemble the data held by other clients. This involves training a generator $G$ with local non-IID data to bridge the gap between datasets. The GAN architecture basically consists of a generator $G$ and a discriminator $D$. In our case, the generator $G$ comprises a series of ’deconvolution’ layers that progressively transform random noise into a sample while the discriminator $D$ closely resembles the global model, except for its output layer, which distinguishes between fake and real samples. The attacker iteratively trains the generator $G$ locally with the constraint of discriminator $D$ until it converges to generate realistic fake samples that do not belong to the attacker. Concurrently, as the SFL procedure progresses, the global model tends to converge. During each server-client communication round, the attacker updates the discriminator $D$ using the latest global model $w_{g}$ downloaded from the server and performs new-round optimization training on the generator $G$, guiding it to generate more authentic fake samples that closely resemble data from other clients. These high-quality, synthetic samples are then integrated into the attacker’s original non-IID dataset, effectively supplementing it with additional data classes.
The attacker synchronously updates the discriminator $D$ during each local training round using the new global model $w_{g}$ downloaded from the server, followed by GAN training to optimize the generator $G$ for improved performance. The output of this process is then merged into the attacker’s non-IID dataset for further backdoor training. As the generated samples closely resemble those from other client datasets, the attacker local optima trained by the attacker can converge closer to the global optima than other clients. Figure 3 demonstrates the difference between the aggregated global optima with and without data supplementation techniques.
Figure 3: Left: global optima under non-IID scenario. Right: global optima with data supplementation techniques.
Trigger selection and injection. With the prepared dataset $D_{p}$, containing both original and synthetic data, the attacker proceeds to inject a backdoor into the model. BadSFL leverages three techniques to inject backdoors: (1) Label-flipping (Li et al. 2021a), in which the ground truth labels of a whole class $a$ in $D_{p}$ are directly altered to another label $b$. For instance, all the ’dog’ labels are altered to ’cat’ in CIFAR-10. (2) Pattern trigger (Uprety and Rawat 2021), which involves poisoning samples with a specific trigger pattern, i.e., a small mosaic cube added in the images to activate the backdoor behavior. The attacker injects these poisoned images into the $D_{p}$ along with a target label, establishing a correlation between the trigger pattern and the desired misclassification. (3) A stealthier backdoor method, known as feature-based backdoor (Mayerhofer and Mayer 2022), involves selecting a distinctive feature within an image class as the backdoor trigger. This approach eliminates the need to directly manipulate the images, thereby increasing the difficulty of detection. For instance, all the green cars in the ćarćlass in CIFAR-10 as the backdoor trigger. During the inference stage, the compromised model outputs the attacker’s target label only when the input is an image containing a green car. The selection of a unique feature within a class makes this trigger difficult to detect as it appears as a natural variation within the data.
Backdoor training with control variate optimizing. As discussed in section Challenges with Backdoor Attacks in SFL, the global control variate $c$ is utilized in SFL to correct the client drift. Specifically, the correction value $c-c_{i}$ adjusts the local model point towards the global model, as shown in algorithm 2 in the Appendix. During the local model training process, this correction term effectively ’drag’ the drifting local model closer to the global model, facilitating convergence towards the global optima, as depicted in Figure 4. In the server aggregation round, the global control variate $c$ is computed by averaging the drift values of all local models, which represents the convergence direction of the global model.
Figure 4: Scaffold correction term on a single client.
From the attacker’s perspective, allowing the control variate to correct the poisoned model according to SFL rules can reduce the effectiveness of the backdoor attack, as discussed in Section Challenges with Backdoor Attacks in SFL. However, the attacker still needs to submit a control variate $c_{p}$ to report the drift of the backdoor model. The key idea is to train a backdoor model that is closer to the global model compared to other local models trained on non-IID data. Since the global control variate $c$ is known to participating clients, it can be used as a reference for the global model’s convergence direction, helping to align the poisoned model more closely with the global optimum. This constraint, derived from $c$, functions similarly to using the future global model for optimization, as suggested by Wen et al.(Wen et al. 2022). This constraint can be integrated into the loss function to enhance the backdoor’s effectiveness and persistence in the global model.
Initially, the attacker performs backdoor training and optimizes their backdoor objective as in the Equation 1 (Bhagoji et al. 2019):
$w^{*}_{p}=\mathop{argmin}_{w_{p}}L(D_{p},w_{p}).$ (1)
where $L$ is the loss function of the backdoor task, $w_{p}$ is the attacker model weights.
In our BadSFL attack, we modify the standard backdoor objective function by adding a term to ensure that the backdoor updates sent to the server persist in the backdoor function in the global model for more future training rounds. We can simulate an aggregation round and apply the control variate $c$ to obtain a predicted global model for one future round. Here is the modified objective function (Equation 2):
$P_{j}(w_{p},c)=\frac{w_{p}+w_{g}*(n-1)}{n}-\eta_{l}*c*j$ (2)
To summarize, we formalize our backdoor objective as in Equation 3:
$w^{*}_{p}=\mathop{argmin}_{w_{p}}[L(D_{p},w_{p})+L(D_{p},P_{j}(w_{p},c))].$ (3)
where j represents the number of future rounds that $w_{p}$ anticipates. By optimizing the backdoor model to be closer to the global model, the attacker simultaneously optimizes the control variate $c_{p}$ to align it with the expected drift value. This ensures that the attacker’s actions conform to the SFL protocol 2.
Evaluation
Setup
We evaluate the performance of BadSFL on a server running Ubuntu 18.04 with an NVIDIA GeForce RTX 2080 Ti GPU.
Datasets, models, and hyperparameters. We consider the datasets that are commonly used in previous works (Lyu et al. 2023), i.e., MNIST (Deng 2012), CIFAR-10 (Krizhevsky, Hinton et al. 2009) and CIFAR-100 (Krizhevsky, Hinton et al. 2009). The datasets, hyperparameters, and model structures are summarized in Table 1.
Table 1: Dataset, model structure, and hyperparameter description.
Dataset Instances Features Model Benign $l_{r}$ E Poison $l_{r}$ Poison ratio Batch Size
FL PT FB
CIFAR-10 60000 1024 ResNet-18 0.001 10 0.05 0.1 0.0125 0.01 128
CIFAR-100 60000 1024 ResNet-50 0.0001 100 0.0001 0.1 0.0125 0.01 128
MNIST 70000 784 ConvNet 0.01 2 0.001 0.1 0.0125 - 128
GAN models. In the case of attacking SFL with the CIFAR-10 and CIFAR-100 datasets, the discriminator structure closely mimics the ResNet18 and ResNet50, respectively, with only output layer difference. The corresponding generator comprises five deconvolutional layers that contribute to the generation of synthetic images that can supplement the dataset $D_{p}$ for the backdoor attack. For models on MNIST, the structure of the discriminator closely mimics the LeNet-5. Additionally, the corresponding generator contains three deconvolutional layers strategically configured to generate synthetic images resembling those held by other clients.
Attack bsaselines. We compare BadSFL with two existing backdoor attacks: (1) Black-box Attack (Bagdasaryan et al. 2020): This attack directly poisoned the training dataset through techniques such as label-flipping, pattern trigger and feature-based trigger. Subsequently, it conducts backdoor training to minimize the classification loss on the dataset $D_{p}$. (2) Neurotoxin (Zhang et al. 2022): In addition to the basic attacks, Neurotoxin incorporates a strategic approach to ensure the durability of the backdoor function within the global model. Specifically, it aims to prevent the malicious model updates from pointing towards coordinates that are frequently updated by benign clients, thereby mitigating the risk of the backdoor being erased. Further details and illustrative figures are provided in Appendix8.
Attack settings. We conduct 100 communication rounds in SFL for CIFAR-10, CIFAR-100, and MNIST. Within this framework, 20 clients actively participated in the federated learning system, with each round randomly selecting 50% of them to contribute to the training process. To create a non-IID data distribution, we divided the training dataset into 200 groups sorted by labels and randomly assigned these groups to all the participating clients.
Evaluating our BadSFL attack on the CIFAR-10 and CIFAR-100 datasets, the attacker strategically joins the training process from the 10th round, exiting after the 40th round. Initially, the attacker conducts GAN training to perform data supplementation, employing 10 local epochs, a fixed learning rate of 0.001, and Adam Optimizer with parameters (0.5, 0.999). Subsequently, backdoor training and optimization are carried out with 10 local epochs, a future round $j$ set to 10, a fixed learning rate of 0.05, and SGD optimizer with a momentum of 0.9 and a weight decay of 0.005.
To poison the dataset for obtaining $D_{p}$, we employ three types of backdoor trigger injections and evaluate the performance of BadSFL based on backdoor task accuracy in each communication round, contrasting it with baseline attacks: (1) Label Flipping. All ”dog” samples in $D_{c}$ are relabeled as ”bird”; (2) Pattern Trigger. A small triangle pattern is added to the right bottom of the image, with all the modified samples labeled as ”cat”; (2) Feature-based trigger. Characteristic features are chosen as backdoor triggers in Dataset $D_{c}$, with the corresponding samples labeled as ”bird”. Various features, such as car stripe, green car, red car, plane in sunset, white horse, red ship and yellow truck are selected as the target features for the backdoor injection.
For MNIST, the attacker engages in the training process from the 10th round and quits after the 40th round. Also, GAN training is employed for data supplementation, comprising 2 local epochs, a fixed learning rate of 0.001, and Adam Optimizer with parameters (0.5, 0.999). Subsequently, backdoor training and optimization are performed over 2 local epochs, considering future rounds $j$ as 10, with a fixed learning rate of 0.1 and SGD optimizer.
To poison the dataset for obtaining $D_{p}$, only two types of backdoor trigger injections are utilized, as the information contained in the MNIST dataset is insufficient to identify characteristic features for feature-based backdoor attacks. The performance of BadSFL is evaluated based on the backdoor task accuracy in each communication round, comparing to two baseline attacks: (1) Label flipping. All samples labeled as “5” in $D_{c}$ are relabeled as “2”; (2) Pattern trigger. A small triangle pattern is added to the right bottom of the image, and the labels of all modified samples are changed to “2”.
Evaluation metric. We adopt two established metrics commonly utilized in FL scenarios. The first metric is Primary Task Accuracy (PTA), which reflects the accuracy of the primary classification task on clean samples. The second metric is backdoor task accuracy (BTA), which benchmarks the success rate of attack by measuring the classification accuracy of the poisoned model on poisoned samples.
Results of GAN
Figure 5(a) and Figure 5(c) illustrate the datasets $D_{f}$ generated by the Generator $G$ for CIFAR-10 and MNIST, respectively. In both cases, $D_{f}$ encompasses nearly all classes of each dataset, demonstrating the successful acquisition of full dataset distribution knowledge within the SFL framework under non-IID scenarios. Utilizing the combined dataset $D_{c}$, the attacker executes the backdoor attack while maintaining the model’s accuracy on the primary task. For CIFAR-10, as shown in Figure 5(b), the primary task accuracy remains around 55% with data supplementation, compared to a drop below 25% in the baseline attack. Similarly, for MNIST, Figure 5(d) reveals that using $D_{c}$ instead of $D_{i}$, the primary task accuracy stays above 90%, while the baseline attack results in a decline to under 75%.
Effectiveness of BadSFL
From Figure 6(a) to 6(f), we present a selection of results obtained on the CIFAR-10 and CIFAR-100 datasets (More results available in Figure 7). It is evident that BadSFL outperforms the baseline attacks in terms of both effectiveness and durability. To be specific, BadSFL achieves above 80% backdoor task accuracy across all types of backdoor attacks within the first 10 rounds, while the attacker remains active in the training process for backdoor training and malicious updates to the server. Meanwhile, BadSFL keeps the primary task accuracy at 60% (5(b)). Furthermore, even after the attacker exits the training process at the 40th round, the benign clients continue submitting normal updates in subsequent rounds, which could potentially affect the poisoned updates from the attacker in previous attacking rounds thus erasing the backdoor function. Despite this, BadSFL ensures a resilient backdoor function with accuracy exceeding 90% over the entire 100 SFL rounds, which is 3 times longer than the lifespan achieved by the two baseline attacks, where backdoor task accuracy drops below 50% after the 60th round. Horizontally comparing the effects of different types of backdoor trigger injections (Figure6(a) Figure6(b) Figure6(c)), it is observed the feature-based trigger performs the best among them, benefiting from its stealthiness without directly manipulating the images, thereby making its updates less likely to conflict with those from benign updates.
Figure 6(g) and 6(h) showcases the results obtained on the MNIST dataset. Similarly, BadSFL outperforms the two baseline attacks, achieving both backdoor task accuracy and primary task accuracy above 85%. After the malicious update injection stops at round 40, in the Label flipping attack, the backdoor task accuracy in two baseline attacks catastrophically drops below 40% within 10 rounds whereas BadSFL maintains a 5 times longer-lasting backdoor function in the global model in future rounds. In the pattern trigger attack, BadSFL also injects a more effective backdoor function into global model with 10% higher accuracy comparing to the baselines.
(a) Attacker’s GAN Generated Samples
(b) Primary task accuracy under basic attack and BadSFL attack
(c) Attacker’s GAN Generated Samples
(d) Primary task accuracy under basic attack and BadSFL attack
Figure 5: Dataset Supplementation on CIFAR-10 and MNIST.
(a) Label Flipping
(b) Pattern Trigger
(c) Plane in Sunset
(d) Label Flipping
(e) Plane in Sunset
(f) Yellow Truck
(g) Label Flipping
(h) Pattern Trigger
Figure 6: Backdoor attacks on CIFAR-10 (a, b, c), CIFAR-100 (d, e, f), and MNIST (g, h).
(a) Green Car
(b) Red Car
(c) White Horse
(d) Red Ship
Figure 7: Feature-based backdoor attacks on CIFAR-10.
(a) Label Flipping (CIFAR-10)
(b) Pattern Trigger (CIFAR-10)
(c) Feature-based Trigger
(CIFAR-10)
(d) Feature-based Trigger
(CIFAR-100)
Figure 8: Defense against backdoor attacks.
Defense. Different from previous works, we further conduct multiple defense strategies to demonstrate the superiority of BadSFL. Specifically, we use differential privacy (Xie et al. 2021), model pruning (Liu, Dolan-Gavitt, and Garg 2018), and anomaly detection (Cao et al. 2020) to resist backdoor attacks. Figure 8 shows that BadSFL can still maintain high effectiveness under multiple defense strategies. Taking the label flipping on CIFAR-10 as an example, it can be observed that our method persistently maintains a high attack success rate even under defense strategies, while other baselines fail to attack the model.
Analysis on Neurotoxin. Neurotoxin aims to address the durability issue of backdoor attacks in FL settings, which ensures that the backdoor persists in the global model even after the attacker stops uploading poisoned updates. During FL training, Neurotoxin leverages the concept of the L2 norm, a mathematical function that represents the magnitude of a vector. Neurotoxin observes that the majority of the L2 norm of the aggregated benign gradient is contained in a small number of coordinates, which implies that benign updates tend to cluster in a narrow range and consequently, the aggregated gradient direction is likely to point towards this cluster. Therefore, Neurotoxin identifies and targets the parameters that get minimal changes in magnitude during training, as illustrated in Figure 9 in the Appendix. These relatively stable parameters are less likely to be significantly affected by benign updates, thereby mitigating the risk of the backdoor being overwritten by future updates.
However, our experimental results shown above do not demonstrate the anticipated effectiveness of Neurotoxin. Despite employing this attack strategy, the backdoor accuracy exhibits a similar decline as the baseline attack after the attacker exits the training process at round 50. This observation is also present in our BadSFL experiments and needs further investigation.
Conclusion
This paper introduces BadSFL, a novel backdoor attack specifically tailored for non-IID federated learning environments utilizing the Scaffold aggregation algorithm. By employing a GAN-based data augmentation technique and exploiting the Scaffold’s control variate, BadSFL achieves superior effectiveness, stealthiness, and durability compared to existing methods. Our experimental results on multiple benchmark datasets demonstrate the attack’s effectiveness, with the backdoor persisting significantly longer than existing approaches. In the future, we hope researchers can design more robust defense mechanisms to safeguard federated learning systems against such attacks.